# The brightest X-ray AGNs at redshift $3\lesssim z\lesssim 6$

### v1

@@ 5 Results
The main result of this work is shown in Fig. 4. The figure shows the profile of the extreme-value pdfs (showing the peak values in black, plus the 95th and 99th percentile bands) across redshift bins in the range $3\leq z\leq 6$, assuming five different luminosity functions discussed in Section 2. The step-like feature comes from the fact that we calculate a single EVS distribution in each redshift bin of width 0.5. We include data points from the S82X catalogue described in Section 4.2, and also show the Eddington limit for a selection of these AGN as reported in LaMassa et al. (2024).
Overall we see general consistency between the brightest S82X data and the EVS predictions, which are around 1-2 of magnitudes in $L_{X}$ below the Eddington limits. We also note the following.
  • We see three AGNs that seem to be in slight tension with the EVS predictions of extreme luminosities. These are 2CXO J001217.1-005437 and 2CXO J021043.1-001817 from Chandra, and 4XMM J000748.9+004119 from XMM-Newton, with redshifts $z\approx 3.60,4.73,4.83$. respectively. For certain luminosity functions (Georgakakis et al. (2015), Vito et al. (2014), Ueda et al. (2014)), these AGNs lie above the expected peak values.
  • The Vito et al. (2014) model has the narrowest EVS bands, centering around lower values of $L_{X}$ than other models (i.e. the values of the EVS parameters $\alpha$ and $\beta$ in Table 2 are the smallest amongst the models). The brightest AGNs straddle the outer edges of this band, showing some tension. The hard X-ray AGN luminosity function in Vito et al. (2014) may be biased due to its calibration from soft-band X-ray AGN samples and assuming a simple power-law spectrum with photon index $\Gamma=1.8$. This could explain the sharp drop in $L_{\rm X}$ at high luminosities in Fig. 1 in contrast with other luminosity functions.
@@
  • The Pouliasis et al. (2024) (PDE) and the Ueda et al. (2014) models have similar profiles, with the PDE model predicting slightly higher values for extreme luminosity over the redshift range.
[…]
@@ 6 Discussion and conclusion
In summary, using the EVS formalism, we obtained the prediction for the extreme X-ray luminosities of AGNs in the redshift range $3\lesssim z\lesssim 6$. We derived the probability distribution of extreme luminosities, which we found to be well approximated by the Gumbel distribution. Our main results are shown in Fig. 4 and 5, where we compare observational data with our EVS predictions assuming 5 different semi-analytic models of the luminosity functions. We find good agreement between the models and data, with the Vito model showing most tension with data. The upper end of the extreme luminosities are also found to be generally consistent with Eddington luminosities, but arguably in some tension with the Pouliasis models.
@@
  • At lower redshifts: Elías-Chávez et al. (2024) analyzed the X-ray properties of 23 brightest AGNs from the XMM-Newton Ultra Narrow Deep Field survey, with redshifts up to 2.66. Their 2–10 keV luminosity was found to be $L_{\rm X}\sim 10^{42}$$10^{46}$ erg s-1. Despite ranking among the most luminous AGNs observed in X-ray surveys to date, their luminosities are still consistent with the EVS predictions for all luminosity functions investigated here when extrapolated at the lower end.
@@
    Figure 5: A comparison of EVS predictions for various luminosity functions from Section 2, extended to redshift $z=11$. We include data points from S82-XL for $z>3$, as well as UHZ-1 (Goulding et al., 2023) at redshift $z=10.1$. Each dashed line shows the peak of the extreme-value pdf for each luminosity function (percentile bands not shown).
    In Fig. 5, we plot the peak values of the extreme-value pdf, smoothed over redshifts, extended to $z=11$. These lines correspond to the smoothed version of each model shown in Fig. 4 up to $z=6$. We see that at $z\sim 10$, the peak of the EVS distribution is around $10^{44}-10^{45}$ erg s-1 for all models considered in this paper.
@@ Acknowledgements
C.H. is supported by the Warwick Mathematics Institute Centre for Doctoral Training, and gratefully acknowledges funding from the University of Warwick and the UK Engineering and Physical Sciences Research Council (grant number: EP/W524645/1). P.C. acknowledges support from (i) Suranaree University of Technology (SUT), (ii) Thailand Science Research and Innovation (TSRI), and (iii) National Science, Research and Innovation Fund (grant number 204265).
@@
Appendix A

### v2

@@ 5 Results
The main result of this work is shown in Fig. 4. The figure shows the profile of the extreme-value pdfs (showing the peak values in black, plus the 95th and 99th percentile bands) across redshift bins in the range $3\leq z\leq 6$, assuming five different luminosity functions discussed in Section 2. The step-like feature comes from the fact that we calculate a single EVS distribution in each redshift bin of width 0.5. We include data points from the S82X catalogue described in Section 4.2. The selection criteria described in the previous section explain why our sample size in Fig. 4 appears to be smaller than those reported in Peca et al. (2024).
Overall we see general consistency between the brightest S82X data and the EVS predictions, and note the following.
  • We see three AGNs that seem to be in slight tension with the EVS predictions of extreme luminosities. These are 2CXO J001217.1-005437 and 2CXO J021043.1-001817 from Chandra, and 4XMM J000748.9+004119 from XMM-Newton, with redshifts $z\approx 3.60,4.73,4.83$. respectively. For certain luminosity functions (Georgakakis et al. (2015), Vito et al. (2014), Ueda et al. (2014), Peca et al. (2023)), these AGNs lie above the expected peak values.
  • The Vito et al. (2014) model has the narrowest EVS bands, centering around lower values of $L_{\rm X}$ than other models (i.e. the values of the EVS parameters $\alpha$ and $\beta$ in Table 2 are the smallest amongst the models). The brightest AGNs straddle the outer edges of this band, showing some tension. The hard X-ray AGN luminosity function in Vito et al. (2014) may be biased due to its calibration from soft-band X-ray AGN samples and assuming a simple power-law spectrum with photon index $\Gamma=1.8$. This could explain the sharp drop in $L_{\rm X}$ at high luminosities in Fig. 1 in contrast with other luminosity functions.
@@
  • The Pouliasis et al. (2024) (PDE), the Ueda et al. (2014) and the Peca et al. (2023) models have similar profiles, with the PDE model predicting slightly higher values for extreme luminosity over the redshift range.
@@ 6 Discussion and conclusion
Using the EVS formalism, we obtained the prediction for the extreme X-ray luminosities of AGNs in the redshift range $3\lesssim z\lesssim 6$. We derived the probability distribution of extreme luminosities, which we found to be well approximated by the Gumbel distribution. Our main results are shown in Fig. 4 and 5, where we compare observational data with our EVS predictions assuming 5 different semi-analytic models of the luminosity functions. We find good agreement between the models and data, with the Vito model showing most tension with data.
While there might be some biases from survey sensitivity, completeness, and absorption distribution, these factors are largely not critical to our specific analysis. For the brightest AGNs whose fluxes are well above the S82X flux limits, biases related to absorption and survey sensitivity have minimal impact on the interpretation of our EVS results. Future studies could test this assumption, for example, by using absorption-corrected luminosities where available. Nevertheless, our study does not aim to model the entire underlying AGN population, nor do we attempt to constrain the full AGN XLF. Instead, our focus is on comparing the brightest, confirmed AGN observed in S82X to the upper envelope of AGN luminosities predicted by extreme value statistics. One of our key results is that the brightest X-ray luminosities in S82X mostly fall below the theoretical maximum predicted by EVS assuming various LF models. This conclusion holds true regardless of the number of undetected or excluded fainter AGNs.
@@
  • At lower redshifts: Elías-Chávez et al. (2024) analysed the X-ray properties of 23 brightest AGNs from the XMM-Newton Ultra Narrow Deep Field survey, with redshifts up to 2.66. Their 2–10 keV luminosity was found to be $L_{\rm X}\sim 10^{42}$$10^{46}$ erg s-1. Despite ranking among the most luminous AGNs observed in X-ray surveys to date, their luminosities are still consistent with the EVS predictions for all luminosity functions investigated here when extrapolated at the lower end.
@@
    Figure 5: A comparison of EVS predictions for various luminosity functions from Section 2, extended to redshift $z=11$. We include data points from S82-XL for $z>3$, as well as UHZ-1 (Goulding et al., 2023) at redshift $z=10.1$. Each line shows the peak of the extreme-value pdf for each luminosity function (percentile bands not shown).
    In Fig. 5, we plot the peak values of the extreme-value pdf, smoothed over redshifts, extended to $z=11$. We have shown the redshift range $3<z<11$, to avoid overcrowding of the data points in the low redshift range $3<z<3.5$. The lines correspond to the smoothed version of each model shown in Fig. 4 up to $z=6$. We see that at $z\sim 10$, the peak of the EVS distribution is around $10^{44}-10^{45}$ erg s-1 for most models considered in this paper with the exception of the Peca et al. (2023) model, which is most likely due to extrapolating too far out of the redshift range used to the construction of the numerical fit.
@@ Acknowledgements
We would like to thank to Bovornpratch Vijarnwannaluk and Alessandro Peca for their helpful suggestions. C.H. is supported by the Warwick Mathematics Institute Centre for Doctoral Training, and gratefully acknowledges funding from the University of Warwick and the UK Engineering and Physical Sciences Research Council (grant number: EP/W524645/1). P.C. thanks funding support from the National Science Research and Innovation Fund (NSRF) via the Program Management Unit for Human Resources & Institutional Development, Research and Innovation [grant number B46G670083]).
@@
Appendix A Uncertainties in Number Count
In this section we discuss the uncertainties that could arise in the number count in equation (9). We incorporate the possible uncertainties into one constant $f_{\rm unc}$ defined through the following equation:
$N$ $=$ $f_{\rm unc}\iiint f_{\rm abs}(\log L_{\rm X},z,N_{\rm H})\frac{\,\textrm{d}\Phi(\log L_{\rm X},z)}{\,\textrm{d}\log L_{\rm X}}$ (16)
$\times f_{\rm sky}\frac{\,\textrm{d}V}{\,\textrm{d}z}\,\textrm{d}\log L_{\rm X}\ \,\textrm{d}z\ \,\textrm{d}\log N_{\rm H}.$
In Fig. 6 we plot the EVS predictions for $f_{\rm unc}=0.5,0.8,1,1.2,1.5$, where $f_{\rm unc}=1$ corresponds to no uncertainties in the number count. From this figure, we see that even for the large uncertainties, the peak luminosity changes relatively little ($\sim\pm 20\%$).
We expect uncertainties from redshift incompleteness (with $f_{\rm unc}<1$) to have negligible impact on our analysis. For instance, the $S82X$ catalogue of Peca et al. (2023) has almost $100\%$ redshift completeness, and for our subset of data points in Fig. 4, we only included those with spectroscopically confirmed redshifts.
We used $f_{\rm unc}>1$ to illustrate the opposite effects from various factors such as Malmquist bias (describing the loss of faint sources due to the flux limit in a survey) or uncertainty in the value of $f_{\mbox{\scriptsize{sky}}}$. Both Vito et al. (2014) and Georgakakis et al. (2015) include an uncertainty factor as a function of $(L_{\rm X},\ z,\ N_{\rm H})$ to account for the redshift incompleteness in the AGN samples they use.
While these uncertainties are important to acknowledge, we can see here that they do not meaningfully change our EVS calculations of peak luminosities.
Figure 6: A comparison of EVS predictions for the Pouliasis et al. (2024) PDE luminosity model when applying uncertainty factors of $f_{\rm unc}=0.5,0.8,1,1.2,1.5$, where the black dashed line shows the peak luminosity with no uncertainty, or $f_{\rm unc}=1$, applied to the number count defined in equation (16). This curve corresponds to the smoothed EVS profile shown in Fig. 4.
Appendix B GEV parameters